OpenStudy (anonymous):

need help finding the intercept x and y for y=(-3/4)x+5

OpenStudy (anonymous):

anybody out there?

OpenStudy (anonymous):

compare with y=mx+c m=gradient c=y intercept

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y intercept=5

OpenStudy (anonymous):

and for intercept x put y=0

OpenStudy (anonymous):

If you want the x intercept, set y = 0 if you want the y intercept, set x = 0

OpenStudy (anonymous):

in this case its better to compare it to avoid working !!!!

OpenStudy (anonymous):

yes. i did that but i think im doing the math wrong.

OpenStudy (anonymous):

WHAT!

OpenStudy (anonymous):

got a scanner can show you what i've done. just really need help!

OpenStudy (anonymous):

OK !!

OpenStudy (anonymous):

would you like to see?

OpenStudy (anonymous):



OpenStudy (anonymous):

what r u doing ????

OpenStudy (anonymous):

hold on.

OpenStudy (anonymous):

0=-3/4x+5 , -5=-3/4x, -20=3x, ??? just confused..please help?

OpenStudy (anonymous):